OpenStudy (anonymous):

find a quadratic equation with integer coefficients with no common factors, which has the given numbers as solutions: x=3, x=3/4 ___=0

OpenStudy (anonymous):

(x-3)*(x-3/4)=0 (x-3)*(4x-3)=0 \[4x^2-15x+9=0\]